OpenStudy (anonymous):

Can someone PLEASE show me how to do 2√u(√5 + 3√45) ?

OpenStudy (mertsj):

First simplify sqrt45

OpenStudy (mertsj):

What did you get?

OpenStudy (anonymous):

√9√5 ?

OpenStudy (mertsj):

And what is sqrt9?

OpenStudy (anonymous):

3

OpenStudy (mertsj):

So you need to multiply it by the coefficient 3 and get 9sqrt5

OpenStudy (mertsj):

Now inside the parentheses you have similar radicals so add them together.

OpenStudy (mertsj):

What do you get?

OpenStudy (anonymous):

Wait so now would the expression look like this?... 2√u ( √5 + 3√5) ?

OpenStudy (mertsj):

No. What happened to the 3?

OpenStudy (anonymous):

It turned to a 9 ah ok so then

OpenStudy (anonymous):

2√u (10√5)

OpenStudy (mertsj):

\[3\sqrt{45}=3\sqrt{(9)(5)}=3(3)\sqrt{5}=9\sqrt{5}\]

OpenStudy (mertsj):

Yes now multiply the 2 and 10 and then the radicands.

OpenStudy (anonymous):

20√5u, hey that's one of the answers

OpenStudy (mertsj):

Cool Good job.

OpenStudy (anonymous):

Thanks so much you actually helped me through it

OpenStudy (mertsj):

yw